# Be like a Goldfish, Don’t Memorize! Mitigating Memorization in Generative LLMs

### v1

@@ Abstract
Large language models can memorize and repeat their training data, causing privacy and copyright risks. To mitigate memorization, we introduce a subtle modification to the next-token training objective that we call the goldfish loss. During training, a randomly sampled subset of tokens are excluded from the loss computation. These dropped tokens are not memorized by the model, which prevents verbatim reproduction of a complete chain of tokens from the training set. We run extensive experiments training billion-scale Llama-2 models, both pre-trained and trained from scratch, and demonstrate significant reductions in extractable memorization with little to no impact on downstream benchmarks.
@@ 1 Introduction
We present the goldfish loss, a strikingly simple technique that leverages properties of the next-token prediction objective to mitigate verbatim generation of memorized training data. Like standard training, the proposed approach begins with a forward pass on all tokens in a batch. Unlike standard training, in which the next token prediction loss is calculated on all inputs, we exclude a pseudo-random subset (e.g., 25%) of the training tokens. On the backward pass, the model never learns to reproduce the excluded tokens. At inference time, the model must make a “guess” each time it runs into a dropped token, causing it to depart from the training data sequence. In this way, the goldfish loss enables training on text without the ability to make a verbatim reproduction at inference time.
Our exploration of this idea begins by stress-testing the goldfish loss with a training setup that aggressively promotes memorization. We train a 7B parameter model on a small number of articles for 100 epochs, finding that the models trained with goldfish loss resist memorization while standard training memorizes most of the training data (See Figure 1). We then turn to more standard training regimens, where we observe that the memorization metrics of goldfish models closely resemble models that never saw the training data at all. We then look at the utility of goldfish models and observe that they still learn effectively from training data, although in some situations they may need to train for longer than standard models to compensate for the lost tokens that were excluded from the loss. Finally, we try to adversarially extract training data from goldfish models using an aggressive beam search decoder, but this typically fails. We do, however, observe that membership inference attacks still work on goldfish models, albeit with marginally lower accuracy.
@@ 2.1 Quantifying Memorization in LLMs
Both benign and adversarial prompting strategies can both extract training data from open-sourced large language models (Carlini et al., 2019, 2021; Inan et al., 2021). Carlini et al. (2023) proposes a family of concrete memorization metrics including “extractable memorization” with prefix length $k$, where if the model memorizes a string, it will regurgitate the rest of the string when prompted with a prefix of length $k$. This notion of memorization is the focus of the work below, as it represents a worst-case scenario and is easy to reproduce in controlled experiments. It should be noted that training data can be extracted without using a $k$-prefix. Spontaneous reproducing of training data has been observed in both language models (Nasr et al., 2023) and image generators (Somepalli et al., 2023) without any prior knowledge of the data content. More recently, Schwarzschild et al. (2024) proposes a novel definition for memorization that quantifies whether a training string is extractable by an adversarial prompt that is shorter than the string itself.
@@ 2.2 Mitigating Memorization in LLMs
It is known that deduplicating training data can mitigate memorization (Kandpal et al., 2022). However, this is complicated by the scale of web data and the prevalence of near-duplicated versions of many texts. Distinct from work on training time techniques, Ippolito et al. (2022) proposes detection of memorization at test time using a Bloom filter (Bloom, 1970). It should be noted that this approach is also vulnerable to missing near-duplicated documents due to the brittle data structure and feature extractors used.
@@ 2.3 Regularization and Memorization
Classical definitions of memorization relate to overfitting (Feldman and Zhang, 2020) and argue that memorization is reduced through regularization techniques that reduce overfitting, through strategies such as weight decay and dropout (Srivastava et al., 2014). However, both are insufficient to prevent memorization in LLMs (Tirumala et al., 2022; Lee et al., 2022a). Related regularization strategies are the addition of noise to input embeddings (Jain et al., 2024; Wen et al., 2024), or random dropout of tokens during training (Hou et al., 2022). Lin et al. (2024) propose to drop tokens in a data-dependent manner, to train the model only on a subset of tokens, continuously selected by a reference model.
Our approach is conceptually quite different because we forgo randomized regularization, and construct a localized, pseudo-random token mask — every time a certain phrase or passage appears in the data, the passage is masked in the same manner, preventing the model from learning the entire passage verbatim. In comparison, if the model is trained with randomized dropout of tokens or weights, it will eventually learn the entire passage, as the passage is seen multiple times with different information masked.
@@ 3 Goldfish Loss: Learning Without Memorizing
$\mathcal{L}(\theta)=-\frac{1}{L}\sum_{i=1}^{L}\log P(x_{i}|x_{<i};\theta).$ (1)
@@
The goldfish loss is only computed on a subset of the tokens, and thus prevents the model from learning the entire token sequence. Choose a goldfish mask $G\in\{0,1\}^{L}$ and define the goldfish loss as
@@
In plain English, we ignore the loss on the $i$th token if its mask value is $G_{i}=0$, and include the token if $G_{i}=1.$ Most importantly, the outputs $x_{i}$ are still conditioned on all prior tokens $x_{<i}$, allowing the model to learn the full distribution of natural language over the course of training. Yet, for a given passage, the model does not learn to predict the $i$th token, and so is never conditioned on the exact sequence $x_{<i}$ at test time. Note that the goldfish mask will be chosen independently for each training sample, based on local context.
@@
There are a range of ways to choose the goldfish mask, after choosing a drop frequency $k$. A simple baseline that we investigate is to drop every $k$th token in a sequence, which we refer to as a static mask, which we juxtapose with a random mask baseline that drops every token with probability $1/k$. We use the random mask to differentiate the effects of regularization that random dropping provides from the effects of the goldfish loss, which is deterministic.
@@ 4 Can Goldfish Loss Prevent Memorization?
In this section, we validate that the goldfish loss can indeed prevent memorization. We consider two setups, an extreme setup that aggressively promotes memorization with many epochs on few samples, and a standard setup that emulates the kinds of batching used in realistic model training.
We quantify memorization using two metrics. We first chop each test sequence from the training set into a prefix and a suffix of length $n$ tokens. Conditioned on the prefix, we autogressively generate text with zero temperature. We compare the generated suffix with the ground truth suffix using two metrics. These are (1) RougeL score (Lin, 2004) which quantifies the length of the longest common (non-consecutive) subsequence. A score of $1.0$ indicates perfect memorization. (2) Exact Match rate, which measures the percentage of correctly predicted tokens compared to ground truth text.
@@ 4.1 Preventing Memorization in Extreme Scenarios
We begin by considering a training setup that is specifically designed to induce memorization. We train a LLaMA-2-7B model for 100 epochs on a dataset consisting of only 100 English Wikipedia (Wikimedia Foundation, ) articles. We select these documents by randomly sampling a set of pages that contain between $2000$ and $2048$ tokens. In Figure 1, we observe that standard training results in verbatim memorization of $84/100$ articles, while the goldfish loss model with $k=4$ memorized none. RougeL metrics indicate that the model trained with goldfish loss repeats non-consecutive $n$-gram sub-sequences that are roughly twice as long as a model that never saw the data. This is consistent with our definition. The model still memorizes subsequences, but the likelihood of getting a long subsequence correct reduces exponentially in the length of the subsequence.
Figure 3: Benchmark Performance: We pretrain 1B parameter models on 20 billion tokens as described in Section 4.1 and evaluate downstream performance on various benchmarks. We note only marginal change in performance for models trained with goldfish loss (with various values of k) in comparison to the model trained with standard loss. Control refers to model trained only on RedPajama and not on wikipedia documents.
@@ 4.2 Preventing Memorization in Standard Training
Our experimental set-up largely follows that of TinyLLaMA-1.1B (Zhang et al., 2024a). We pretrain a language model of size $1.1$B with a vocabulary size of $32$k. We compare the goldfish loss in Equation 2 at different values of $k$ and the standard causal language modeling loss in Equation 1. More training details can be found in Appendix A.
We construct the dataset for this experiment based on two sources. First, a subset of RedPajama version 2 (Together Computer, 2023), on which we train for a single epoch. Second, we also mix in $2000$ target sequences, each of $1024$ to $2048$ token length, from the Wikipedia (Wikimedia Foundation, ) corpus. To simulate the problematic case of data that is duplicated within the dataset, we repeat this target set $50$ times in the course of training, in random locations. In total, we train on $20$ billion tokens in over $9500$ gradient steps.
@@
Our intuition is that tokens are not memorized when they are dropped by the goldfish loss, leading to model divergence from the ground truth. To validate this intuition, we analyze the relationship between the positions of dropped tokens and the positions at which the model diverges from the ground truth while attempting to regenerate the sequence. We consider the $2000$ documents trained for $50$ epochs in Section 4.2. Figure 4 and Table 1 show the relation between dropped index and diverged index.
@@
The goldfish loss seems to prevent memorization, but what are the impacts on downstream model performance? Does a model still learn effectively? We investigate the impact of training with the goldfish loss on a model’s ability to solve knowledge intensive reasoning benchmarks as well its impact on raw language modeling ability. For most of the downstream evaluations we consider, the knowledge gained from goldfish training is comparable to standard training. When considering language modeling ability, we find that the goldfish loss causes a mild slowdown in pretraining as one would expect from a model that has seen fewer tokens. However, it matches standard pretraining when both are allowed the same number of supervised tokens for loss computation (see Figure 5).
@@ 5.2 Impact on Language Modeling Ability
Because goldfish models have, in a sense, trained on fewer tokens than standard models, we might expect their raw token prediction ability to trail behind standard models that have seen more tokens. We quantify this impact by tracking a model’s token-for-token progress throughout training, as measured by validation loss as well as each model’s ability to complete web-text documents from the training data with high semantic coherence to the ground truth.
@@
To understand the impact on the model’s training progression, we analyze the validation loss in terms of the total number of supervised tokens. In Figure 5, we show the validation loss curves over $12$M tokens of RedpajamaV2 data on three models, one trained with standard loss and another two trained with $4$-GL. The $4$-GL models drop tokens using a pseudo-random hash of width $13$, which is the same setup used in Figure 1.
Figure 5: Validation Loss Curves During Pretraining: We pretrain three models on RedpajamaV2 data, one with standard loss and another two with $4$-GL. Each line indicates the validation loss as a function of supervised tokens (tokens used for loss computation). Goldfish runs need to do a forward pass on more tokens to achieve the same number of supervised tokens. We either increase batch size or add more total steps, bringing the total number of input tokens to $267$B for goldfish runs so that the supervised token count is $200$B for all runs.
Since the net number of supervised tokens is fewer with goldfish loss than with standard loss, we plot the number of supervised tokens (i.e., the tokens used in the loss calculation) against the validation loss of RedPajama-V2. For all models, we train with $200$ billion supervised tokens. This corresponds to $200$ billion input tokens for the standard loss and $267$ billion input tokens for the goldfish loss. The calculation is based on the formula: $(1-\frac{1}{k})\times\texttt{ Input Tokens}=\texttt{Supervised Tokens}$, where $k=4$.
From Figure 5, we observe that all models converge to nearly identical validation loss values. Additionally, both the standard loss and the goldfish loss with increased batch size follow almost the same validation curve. We hypothesize that this is because the total number of supervised tokens per iteration, combined with an aligned learning rate schedule, causes similar progression during training. Moreover, we notice that increasing the total number of steps allows the goldfish loss to advance ahead in training for most of the curve. We suspect this is due to the higher learning rate being maintained for a longer period during training.
We conclude that the goldfish loss performs similarly to the standard loss when both are given the same number of supervised tokens. However, to achieve performance parity, goldfish training requires more tokens to be used on the forward pass to compensate for the tokens ignored in the loss computation.
@@
As an additional confirmation that models trained with goldfish loss retain their ability to produce fluent and faithful outputs, we compute Mauve score (Pillutla et al., 2021), a metric used to evaluate the quality of generated text against real text by measuring similarity in terms of diversity and naturalness.
We present Mauve scores for models trained with goldfish loss on samples from the Slimpajama (Soboleva et al., 2023) dataset in Figure 6. We see that under greedy decoding, there is a minimal drop in Mauve scores as compared to the Control or CLM baseline model under any of the $k$ values tested. However, when generating completions using multinomial sampling with temperature $0.7$, we see scores trend up slightly as $k$ increases and the model sees more tokens. Note that goldfish loss becomes equivalent to the standard CLM objective in the limit of large $k$.

### v2

@@ Abstract
Large language models can memorize and repeat their training data, causing privacy and copyright risks. To mitigate memorization, we introduce a subtle modification to the next-token training objective that we call the goldfish loss. During training, a randomly sampled subsets of tokens are excluded from the loss computation. These dropped tokens are not memorized by the model, which prevents verbatim reproduction of a complete chain of tokens from the training set. We run extensive experiments training billion-scale LLaMA-2 models, both pre-trained and trained from scratch, and demonstrate significant reductions in extractable memorization with little to no impact on downstream benchmarks.
@@ 1 Introduction
We present the goldfish loss, a strikingly simple technique that leverages properties of the next-token prediction objective to mitigate verbatim generation of memorized training data (Section 3). Like standard training, the proposed approach begins with a forward pass on all tokens in a batch. Unlike standard training, in which the next token prediction loss is calculated on all tokens, we exclude a pseudo-random subset (e.g., 25% i.e. with probability $1/4$) of the training tokens. The tokens are dropped with $1/k$ probability where $k$ is a chosen hyperparameter. On the backward pass, the model never learns to reproduce the excluded tokens. At inference time, the model must make an unsupervised “guess” each time it tries to predict a dropped token, causing it to depart from the training data sequence. In this way, the goldfish loss enables training on text without the ability to make a verbatim reproduction at inference time. We formally introduce goldfish loss in Section 3. Throughout the paper, we either use $k=4$ or refer to it as $k$-GL, indicating the value of the drop frequency $k$.
Our exploration of this idea begins by stress-testing the goldfish loss with a training setup that aggressively promotes memorization (Section 4.1). We train a 7B parameter model on a small number of articles for 100 epochs, finding that the models trained with goldfish loss resist memorization while standard training memorizes most of the training data (see Figure 1). We then turn to more standard training regimen, where we observe that the memorization metrics of goldfish models closely resemble models that never saw the training data at all (Section 4.2). We then look at the utility of goldfish models and observe that they still learn effectively from training data (Section 5.1), although in some situations they may need to train for longer than standard models to compensate for the lost tokens that were excluded from the loss (Section 5.2). Finally, we try to adversarially extract training data from goldfish models using an aggressive beam search decoder, which typically fails. We do, however, observe that membership inference attacks still work on goldfish models, albeit with marginally lower accuracy (Section 6).
@@ 2.1 Quantifying Memorization in LLMs
Both benign and adversarial prompting strategies can extract training data from open-sourced large language models (Carlini et al., 2019, 2021; Inan et al., 2021). Carlini et al. (2023) proposes a family of concrete memorization metrics including “extractable memorization” with prefix length $p$, where if the model memorizes a string, it will regurgitate the rest of the string when prompted with a prefix of length $p$. This notion of memorization is the focus of our work, as it represents a worst-case scenario and is easy to reproduce in controlled experiments. It should be noted that training data can be extracted without using a $p$-prefix. Spontaneous reproducing of training data has been observed in both language models (Nasr et al., 2023) and image generators (Somepalli et al., 2023) without any prior knowledge of the data content. More recently, Schwarzschild et al. (2024) proposes a novel definition for memorization that quantifies whether a training string is extractable by an adversarial prompt that is shorter than the string itself.
@@ 2.2 Mitigating Memorization in LLMs
It is known that deduplicating training data can mitigate memorization (Kandpal et al., 2022). However, this is complicated by the scale of web data and the prevalence of near-duplicated versions of many texts. Distinct from work on training time techniques, Ippolito et al. (2022) proposes detection of memorization at test time using a bloom filter (Bloom, 1970) data structure. It should be noted that this approach is also vulnerable to missing near-duplicated documents due to the brittle data structure and feature extractors used.
@@ 2.3 Regularization and Memorization
Classical definitions of memorization relate to overfitting (Feldman and Zhang, 2020) and argue that memorization is reduced through regularization techniques that reduce overfitting, through strategies such as weight decay and dropout (Srivastava et al., 2014). However, both are insufficient to prevent memorization in LLMs (Tirumala et al., 2022; Lee et al., 2022a). Related regularization strategies are the addition of noise to input embeddings (Jain et al., 2024; Wen et al., 2024), or random dropout of tokens during training (Hou et al., 2022). Lin et al. (2024) study dropping tokens from the loss in a data-dependent manner and observe that this can enhance model performance if tokens are carefully selected by a reference model. The idea of dropping parts of each training sample was successfully used to prevent memorization in diffusion models by Daras et al. (2024a, b). Here, images are degraded by removing many patches before they are used for training. While conceptually related to our proposed method, the goldfish loss achieves high efficiency by computing a forward pass on an entire unaltered text sample, and only excluding information from the backward pass.
Our approach is conceptually quite different because we forgo randomized regularization, and construct a localized, pseudo-random token mask — every time a certain phrase or passage appears in the data, the passage is masked in the same manner, preventing the model from learning the entire passage verbatim (details in Section 3.1). In comparison, if the model is trained with randomized dropout of tokens or weights, it will eventually learn the entire passage, as the passage is seen multiple times with different information masked.
@@ 3 Goldfish Loss: Learning Without Memorizing
$\mathcal{L}(\theta)=-\frac{1}{L}\sum_{i=1}^{L}\log P(x_{i}|x_{<i};\theta)$ (1)
@@
The goldfish loss is only computed on a subset of the tokens, and thus prevents the model from learning the entire token sequence. For a choosen a goldfish mask $G\in\{0,1\}^{L}$ and goldfish loss is defined as:
@@
In plain English, we ignore the loss on the $i$th token if its mask value is $G_{i}=0$, and include the token if $G_{i}=1.$ Most importantly, the outputs $x_{i}$ are still conditioned on all prior tokens $x_{<i}$, allowing the model to learn the full distribution of natural language over the course of training. Yet, for a given passage, the model does not learn to predict the $i$th token, and so is never conditioned on the exact sequence $x_{<i}$ at test time. Note that the goldfish mask will be chosen independently for each training sample, based on local context using a hash mask (described in detail in Section 3.1).
@@
There are a range of ways to choose the goldfish mask, after choosing a drop frequency $k$. A simple baseline that we investigate is to drop every $k$th token in a sequence, which we refer to as a static mask, which we juxtapose with a random mask baseline that drops every token with probability $1/k$. We use the random mask to differentiate the effects of regularization that random dropping provides from the effects of the goldfish loss, which is deterministic. For our main results, we use hashed mask which we discuss in next section.
@@ 4 Can Goldfish Loss Prevent Memorization?
In this section, we validate that the goldfish loss can indeed prevent memorization. We consider two setups: an extreme setup that aggressively promotes memorization with many epochs (i.e., repetitions) on a few samples, and a standard setup that emulates the batching used in realistic model training.
We quantify memorization using two metrics. We first chop each test sequence from the training set into a prefix and a suffix of length $n$ tokens. Conditioned on the prefix, we autogressively generate text at zero temperature. We compare the generated suffix with the ground truth suffix using two metrics. These are (1) RougeL score (Lin, 2004) which quantifies the length of the longest common (non-consecutive) subsequence. A score of $1.0$ indicates perfect memorization. (2) Exact Match rate, which measures the percentage of correctly predicted sequences compared to ground truth. Since the focus of our work is syntactical memorization, we focus on these two metrics. The results for semantic memorization (or knowledge retention) can be found in Appendix C.1.
@@ 4.1 Preventing Memorization in Extreme Scenarios
We begin by considering a training setup that is specifically designed to induce memorization. We continue pretraining LLaMA-2-7B model (Touvron et al., 2023) for 100 epochs on a dataset consisting of only 100 English Wikipedia (Wikimedia Foundation, ) articles. We select these documents by randomly sampling a set of pages that contain between $2000$ and $2048$ tokens. In Figure 1, we observe that standard training results in verbatim memorization of $84/100$ articles, while the goldfish loss model with $k=4$ memorized none. RougeL metrics indicate that the model trained with goldfish loss repeats non-consecutive $n$-gram sub-sequences that are roughly twice as long as a model that never saw the data. This is consistent with our definition. The model still memorizes subsequences, but the likelihood of getting a long subsequence correct reduces exponentially in the length of the subsequence.
Figure 3: Benchmark Performance: We pretrain 1B parameter models on 20 billion tokens as described in Section 4.1 and evaluate downstream performance on various benchmarks. We note only marginal change in performance for models trained with goldfish loss ($k=3$ and $k=4$) in comparison to the model trained with standard loss. Control refers to model trained only on RedPajama and not on wikipedia canaries.
@@ 4.2 Preventing Memorization in Standard Training
Our second experimental set-up largely follows that of TinyLLaMA-1.1B (Zhang et al., 2024a). We pretrain a language model of size $1.1$B with a vocabulary size of $32$k. We compare the goldfish loss in Equation 2 at different values of $k$ and the standard causal language modeling loss in Equation 1. More training details can be found in Appendix A.
We construct the dataset for this experiment based on two sources. First, a subset of RedPajama version 2 (Together Computer, 2023), on which we train for a single epoch. Second, we also mix in $2000$ target sequences, each of $1024$ to $2048$ token length, from the Wikipedia (Wikimedia Foundation, ) corpus. To simulate the problematic case of data that is duplicated within the dataset, we repeat this target set $50$ times in the course of training, in random locations. In total, we train on $20$ billion tokens in over $9500$ gradient steps. We also train a corresponding control model that is trained only $20$ billion RedPajama tokens.
@@
Our intuition is that tokens are not memorized when they are dropped by the goldfish loss, leading to model divergence from the ground truth. To validate this intuition, we analyze the relationship between the positions of dropped tokens and the positions at which the model diverges from the ground truth while attempting to regenerate the sequence. We consider the $2000$ documents trained for $50$ epochs in Section 4.2. Figure 4 and Table 1 show the relation between dropped index and first diverged index.
@@
The goldfish loss seems to prevent memorization, but what are the impacts on downstream model performance? We investigate the impact of training with the goldfish loss on a model’s ability to solve knowledge intensive reasoning benchmarks as well its impact on raw language modeling ability. For most of the downstream evaluations we consider, the knowledge gained from goldfish training is comparable to standard training.
@@ 5.2 Impact on Language Modeling Ability
Because goldfish models have, in a sense, trained (or supervised) on fewer tokens than standard models, we might expect their raw token prediction ability to trail behind standard models that have seen more tokens. We quantify this impact by tracking a model’s token-for-token progress throughout training, as measured by validation loss as well as each model’s ability to complete web-text documents from the training data with high semantic coherence to the ground truth.
@@
To understand the impact on the model’s training progression, we analyze the validation loss in terms of the total number of supervised tokens. In Figure 5 (left), we show the validation loss curves over $12$M tokens of RedpajamaV2 data. We find that the goldfish loss causes a mild slowdown in pretraining as one would expect from a model that has seen fewer tokens. However, it matches standard pretraining when both are allowed the same number of supervised tokens for loss computation. Supervised tokens indicate the number of unmasked tokens in the goldfish loss case (affected by the chosen $k$) and are the same as the input tokens for standard loss. As observed in Figure 5 (right), we show nearly identical final validation loss values can be achieved either by training for a longer duration (increasing the number of steps) or by using a larger batch size.
Figure 5: Validation Loss Curves During Pretraining: We measure validation loss on the RedPajamaV2 dataset as training progresses. Left: We observe validation loss as a function of input tokens seen during training. The 4-GL model trail behind the standard loss model for the same number of input tokens. Right: However, when matching the standard loss by the count of supervised tokens—i.e., the number of unmasked tokens—either by increasing the number of steps or by expanding the batch size, we observe a similar final validation loss.
Since the net number of supervised tokens is fewer with goldfish loss than with standard loss, we plot the number of supervised tokens (i.e., the tokens used in the loss calculation) against the validation loss of RedPajamaV2. For all models, we train with $20$ billion supervised tokens. This corresponds to $20$ billion input tokens for the standard loss and $26.7$ billion input tokens for the goldfish loss. The calculation is based on the formula: $(1-\frac{1}{k})\times\texttt{ Input Tokens}=\texttt{Supervised Tokens}$, where $k=4$.
Additionally, both the standard loss and the goldfish loss with increased batch size follow almost the same validation curve. Thus, we recommend that when using $k$-GL, one should use the formula above to appropriately transfer the world batch size from the standard loss run.
We hypothesize that this is because the total number of supervised tokens per iteration, combined with an aligned learning rate schedule, causes similar progression during training. Moreover, we notice that increasing the total number of steps allows the goldfish loss to advance ahead in training for most of the curve. We suspect this is due to the higher learning rate being maintained for a longer period during training (under standard cosine scheduler).
We conclude that the goldfish loss performs similarly to the standard loss when both are given the same number of supervised tokens. However, to achieve performance parity, goldfish training requires more tokens to be used on the forward pass to compensate for the tokens ignored in the loss computation indicating this is not a free lunch.
@@
As an additional confirmation that models trained with goldfish loss retain their ability to produce fluent and faithful outputs, we compute Mauve score (Pillutla et al., 2021), a metric used to evaluate the quality of generated text against real text by measuring similarity in terms of diversity and naturalness. This metric also noted to be highly correlated with human text.
We present Mauve scores for models trained with goldfish loss on samples from the Slimpajama (Soboleva et al., 2023) dataset in Figure 6. We see that under greedy decoding, there is a minimal drop in Mauve scores as compared to the Control or CLM baseline model under any of the $k$ values tested. However, when temperature $0.7$, we see scores trend up slightly as $k$ increases and the model sees more tokens. Note that goldfish loss becomes equivalent to the standard CLM objective in the limit of large $k$.